Find the sum of the infinite geometric series
-12 + 48/5 - 192/25 + ...
here are the answer choices
A) negative 40/3
B)negative 80/3
C) negative 20/3
D) negative 10/3
thank you!

Answer by richwmiller(17219)

S=-12*(1 - (-4/5)^20)/(1 - (-4/5))=-6 2/3
S=t*(1 - r^n)/(1 - r)
r=-4/5
c) -20/3 is -6 2/3